A one-dimensional array of 1279128 noninteracting binary spins, each of magnetic moment 0.074 eV/T, are placed in a 4.9T field. At a temperature of 668.8°K determine (a) The average energy per spin. (b) The entropy of the system. (a)D eV Answer part (a) eVKAnswer part (b) Submit Attempted part (a) 0 times and part (b) 0 times There have been no attempts to answer part (a) There have been no attempts to answer part (b)
Answer #1

a) Average energy per spin is given by

  \bar{E} = \mu B

where,

\mu = 0.074~eV/T \\ \\ B = 4.9~T

hence, we get

\bar{E} = (0.074)(4.9) = 0.363~eV

b) Entropy is given by the relation

S = N \frac{\bar{E}}{T}

where,

N = 1279128 \\ \\ T = 668.8\degree K \\ \\ \bar{E} = 0.363~eV

thus, the entropy is

S= (1279128)\frac{0.363}{668.8} = 694.26~eV/ \degree K